# Non-normal Dynamics on Non-reciprocal Networks: Reactivity and Effective Dimensionality in Neural Circuits

### v1

@@ III.1 Steady state analysis
In the cyclic architecture ($c=\epsilon$), the system’s symmetry ensures that the equations for each node are exactly the same. Then, the equilibrium condition $\bar{x}_{i}=\bar{x}_{j}$ is valid for all couple of indices $i,j$. Moreover, with a reasoning analogous to the one in [13], it is easy to show that also in this case the excitatory and inhibitory activities reach the same steady state $\bar{x}_{i}=\bar{y}_{i}$. Then, as in the decoupled case, the system allows a one-dimensional description.
@@
The generalization of the decoupled system to the cyclic one turns out to be straightforward: it is sufficient to substitute $\gamma_{\mu}$ with $\gamma_{\mu}+\gamma_{\ell}$. Therefore, it follows that $\gamma_{\mu}^{c,\epsilon}=\gamma_{\nu}+\alpha-\gamma_{\ell}$, and the scaling of the non-null solution becomes $\bar{\text{x}}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c,\epsilon}\right)$. Fig. 2(a) and (b) report the bifurcation diagram for the population $x_{1}$ and the scaling of the fixed point around the bifurcation, showing that the curve for the cyclic systems as a function of $\gamma_{\mu}$ coincides with the curve for the decoupled system, once adequately shifted. Moreover, in Fig. 2 we show the scaling of $x_{1}$, as a function of the distance to the bifurcation point. Our numerical analyses reveal that the scaling exponent seems to hold not only for the cyclic and decoupled system, but also for a broad range of values of $c<\epsilon$, i.e. for architectures that are not perfect cycles, that we will refer to as ‘weak-feedback systems’.
This suggests that the cyclic interactions do not fundamentally alter the physics of the system, provided that feedback into $x_{3}$ is preserved. Thus, our bifurcation analysis suggests that the cyclic and decoupled models obey the same physics.
Feed-forward coupling
In the feedforward system, the units are organized in a chain. To fix the ideas, we label the different populations as follows: $x_{3}$ is the initial node of the chain and does not receive any input from other units, $x_{2}$ is in the middle of the chain receiving input from $x_{3}$ and projecting onto $x_{1}$, and $x_{1}$ does not project to any other units, being the final node of the chain (see Fig. 1 a)). The population $x_{3}$ trivially behaves as a decoupled unit, thus becoming active at the bifurcation point $\gamma_{\mu}=\gamma_{\mu}^{c}$. For smaller values of $\gamma_{\mu}$, $x_{2}$ receives no input from $x_{3}$ and therefore also $x_{2}$ stays inactive. The same is true also for $x_{1}$, therefore the whole system is inactive for $\gamma_{\mu}<\gamma_{\mu}^{c}$, while it gets triggered by the activation of $x_{3}$, whose input cascades onto the rest of the chain. More formally, around the bifurcation point, we find:
$\dot{\text{x}_{i}}=$ $(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\alpha\text{x}_{i}-\bar{\mathrm{f}}\text{x}_{i}+\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}+$ (8)
@@
where now $\mathrm{f}$ and its derivatives are no longer calculated at zero but in $\epsilon\gamma_{\ell}\bar{\text{x}}_{i+1}$, where $\bar{\text{x}}_{i+1}$ is an external parameter not affected by $\bar{\text{x}}_{i}$. Let us manipulate Eq. (8), to write it in the form $\dot{\text{x}_{i}}=c+b\text{x}_{i}+a\text{x}_{i}^{2}+\mathcal{O}(x^{3})$, by defining $\delta\equiv\gamma_{\mu}-\gamma_{\mu}^{c}$, $b=\delta$, $c=\frac{\gamma_{\nu}}{\alpha}\delta=\mathcal{O}(1)$ and $a=\frac{\gamma_{\nu}}{\alpha}\delta(\delta+\alpha)^{2}-(\delta+\alpha)=\mathcal{O}(1)$. It follows that for $\gamma_{\mu}>\gamma_{\nu}+\alpha$, Eq. (8) is formally equal to
$\dot{\text{x}_{i}}=(1-\bar{\text{x}}_{i+1})\bar{\mathrm{f}}-\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}^{2}-\frac{1}{2}\bar{\mathrm{f}}^{\prime\prime}(\gamma_{\mu}-\gamma_{\nu})^{2}\text{x}_{i}^{2}+\mathcal{O}(x^{3}),$ (9)
with solutions $\bar{\text{x}}\sim\pm\sqrt{\frac{c}{a}}$. This is the normal form for a saddle-node bifurcation [28].
Furthermore, using $\bar{\text{x}}_{3}=\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)+\mathcal{O}(\gamma_{\mu}^{2})$, we can verify that the bifurcation for $x_{2}$ occurs exactly at $\gamma_{\mu}=\gamma_{\mu}^{c}$, given $\bar{\mathrm{f}}\sim\epsilon\gamma_{\ell}\bar{\text{x}}_{3}$ and moreover
$\bar{\text{x}}_{2}\sim\frac{\sqrt{\epsilon\gamma_{\ell}}}{\alpha}\sqrt{\gamma_{\mu}-\gamma_{\mu}^{c}}.$ (10)
Then, by using Eq. (10), we can obtain the results for $x_{1}$: again we can verify that the bifurcation occurs at $\gamma_{\mu}^{c}$, and the scaling reads
$\bar{\text{x}}_{1}\sim\frac{(\epsilon\gamma_{\ell})^{3/4}}{\alpha}\sqrt[4]{\gamma_{\mu}-\gamma_{\mu}^{c}}$ (11)
Figure 2 (panels (a) and (c)) depicts the bifurcation diagram and the fixed points scaling, in agreement with our analytic findings. When we compare the scaling in the feedforward and cyclic cases, we can conclude that the fixed point of the feed-forward setup is more sensitive to $\gamma_{\mu}$ variations compared to the cyclic setup.
@@ III.2 Non-normal dynamics in non-reciprocal motifs
In this section, first we focus on the calculation of eigenvectors of the Jacobian near the equilibrium points; next, we characterize the non-normality by defining a parameter that quantifies the angle between eigenvectors. This step will be key towards understanding the phenomenology of the stochastic regime. Again, we will focus on highlighting the differences between the cyclic and feedforward architectures.
Non-normal matrices (i.e. matrices that do not commute with their adjoint) typically have non-orthogonal eigenvectors and the angle between such eigenvectors is related to the non-normality of the system: the smaller the angle between two eigenvectors, the more prominent the effects of the non-normality.
Before proceeding with the linear analysis, it is useful to rewrite the system in terms of the following variables:
@@
Since the solution will always be of the form $\text{x}_{i}=\text{y}_{i}$ $\forall i$, the equilibrium coordinate for the second variable will be $\bar{\Delta}_{i}=0$. Then, breaking down $\Sigma_{i}$ and $\Delta_{i}$ into the equilibrium and perturbation components, $\Sigma_{i}=\bar{\Sigma}_{i}+\sigma_{i}$ and $\Delta_{i}=\bar{\Delta}_{i}+\delta_{i}$, we can rewrite the linear expansion of the equations as follows
$\dot{\sigma}_{i}=-\alpha\sigma_{i}-\sigma\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}\biggl{[}\sigma(\gamma_{\mu}-\gamma_{\nu})+\delta(\gamma_{\mu}+\gamma_{\nu})+$ (13)
$\hskip 193.47873pt+\epsilon\gamma_{\ell}\frac{\sigma+\delta}{2}\biggr{]}$
$\dot{\delta}_{i}=-\alpha\delta_{i}-\delta\bar{\mathrm{f}}$
We observe that, in the new coordinate system, the Jacobian associated with the linear dynamics takes an upper triangular form. This change of basis, originally proposed in [13] for symmetry considerations in the decoupled system, effectively implements the Schur transformation of the Jacobian of 5, ensuring its upper triangular structure in the new variables.
The Jacobian matrices are calculated for generic fixed points since we are interested in the derivation of non-normality measures in generic regimes.
Cyclic couplings   Let’s begin with the study of the cyclic system ($c=\epsilon$). The Jacobian of the cyclic model has a block circulant matrix shape and thus it can be block diagonalized according to the circulant matrix theorem [29], as shown in SM. It is interesting to note that the block diagonalization leads to the basis $(\Sigma_{i},\Delta_{i})$. The same result can be achieved by a Schur transformation in the complex field.
A generic circulant matrix is diagonalized by a special unitary matrix known as the discrete Fourier transform matrix, which has $(r,s)$ element equal to $\exp(-2\pi\mathrm{i}(r-1)(s-1)/n)$. In our model, one can write the diagonalized Jacobian for the cyclic $J_{c}^{(D)}=SJ_{c}S^{-1}$ as follows:
@@
$\Phi=\begin{pmatrix}-\alpha-\bar{\mathrm{f}}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\gamma_{\nu})\\ 0&-\alpha-\bar{\mathrm{f}}\end{pmatrix},\hskip 8.53581pt$ (15)
$\psi=\begin{pmatrix}\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}\\ 0&0\end{pmatrix}$
Therefore, according to this base, the graph can be remapped into three disjoint feedforward graphs as depicted in Fig. 1b. Exploiting the upper triangular shape of the sub-matrices, we can read the eigenvalues in the diagonal elements:
@@
where $k_{i}=\biggl{\{}\epsilon,\epsilon\biggl{(}-\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr{)},\epsilon\biggl{(}-\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr{)}\biggr{\}}$, $i=1,2,3$. We identify two complex and four real eigenvalues, wherein global stability is constrained by the inequality $-\alpha-\bar{f}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})<0$. In particular the null fixed point loses its stability when $\gamma_{\mu}=\gamma_{\nu}+\alpha$, as already proved in Sec. III.1.
We are now ready to calculate the form of the eigenvectors. The full derivation, including a generalization to $n$ populations, can be found in SM. From the form of the eigenvectors’ elements, it becomes evident that $\xi=|{\gamma_{\nu}-\gamma_{\mu}-\epsilon\gamma_{\ell}}|/({\gamma_{\mu}+\gamma_{\nu}+\epsilon\gamma_{\ell}})$ controls the angle between two of the six eigenvectors. These two eigenvectors encompass contributions of excitatory and inhihbitory neurons across all populations; in particular one of them is defined by $\Sigma_{i}=1,\Delta_{i}=0$ –i.e. lies in the direction where all excitatory and inhibitory populations have the same firing rate– and the other one is defined by $\Sigma_{i}=1,\Delta_{i}=\xi$.
Thus, by tuning the parameter $\xi$ one can directly modulate the angle between these two eigenvectors, which determines the degree of the non-normality in the system.
In the asymptotic limit $\xi\rightarrow\infty$ the eigenvectors become orthogonal and the system becomes normal. Conversely, when $\xi=0$, $J_{c}^{(D)}$ becomes defective, as the two eigenvectors collapse into one. In this limit, non-normality emerges from a single structural mechanism, controlled exclusively by the parameter $\xi$. In other words, $\xi$ is the sole parameter through which the eigenvector collapse occurs. This parameter $\xi$ can be seen as a generalization of the parameter $\zeta={\gamma_{\nu}-\gamma_{\mu}}/({\gamma_{\mu}+\gamma_{\nu}})$ controlling non-normality in the decoupled system [7]. Furthermore, the degenerate eigenvector defined by $\Sigma_{i}=1,\Delta_{i}=0$ extends the corresponding structure found in the decoupled system.
Once again, by reabsorbing $\epsilon\gamma_{\ell}$ into $\gamma_{\mu}$ the cyclic system maps onto the decoupled one.
Feed-forward couplings   Let’s move on to the feedforward case by setting $c=0$. Now the Jacobian matrix takes the form:
@@
where the subscripts indicate that the 222x222 matrices $\Phi_{k}$ (respectively $\Psi_{k}$) share the same structure, but are evaluated at different equilibrium points, since each unit of the chain reaches a different steady state. However, as stated above, the equilibrium point of the excitatory and inhibitory sub-populations within each unit, remains homogeneous (i.e. $\bar{\text{x}}_{i}=\bar{\text{y}}_{i}$). The upper diagonal form of $J_{f}$ allows to read directly the eigenvalues on the diagonal (see SM), i.e.:
@@
Again, the complete form of the eigenvectors, along with their derivation and a generalization to $n$ populations, are provided in SM.
Interestingly, both the elements of the eigenvectors and the eigenvectors themselves exhibit a structured pattern that reflects a recursive-like organization, explicitly highlighting the directionality of signal propagation along the chain. Moreover, in this case, the non-normality parameter has two contributions. The first, encoding the so-called inner non-normality, is given by $\zeta=|{\gamma_{\nu}-\gamma_{\mu}}|/({\gamma_{\mu}+\gamma_{\nu}})$ which coincides with the non-normality parameter of the decoupled system [7]. The second contribution, representing the outer non-normality, is associated with the directionality of the flow along the chain, and coincides with the coupling parameter $\gamma_{\ell}$.
Similar to the cyclic setup, $J_{f}$ becomes defective for $\zeta\rightarrow 0$ and normal for $\zeta=\infty$. Additionally, in the feedforward system, the non-normality can also be influenced by $\gamma_{\ell}$. In this case, the Jacobian becomes defective also when $\gamma_{\ell}\rightarrow\infty$, while $\gamma_{\ell}$ is not able to alter the inner non-normality.
The parameters that we derived here to tune the angle between the eigenvectors, do not coincide with the Henrici measure, commonly used to quantify the strength of the non-normality [12]. A full comparison is discussed in SM.
Figure 3: Reactivity and comparison of numerical abscissa and non-normality parameters. On the right $\rho(t)$ (panels c), d), e) and f)) trajectories under reactivity. Reactivity increases from left to right with the numerical abscissa $m$. Here $\zeta$ represents the non-normality parameter for the binary system. Top row panels c) and d): In each panel the couplings are chosen for the decoupled system and remain unchanged over the setups. From c) and d) $\xi$ keeps fixed. Bottom row panels e) and f): As the setup changes, the non-normality of both the feedforward and cyclic systems is adjusted to maintain a consistent level across all configurations in each panel. With this change, only the feedforward system exhibits a significant increase in reactivity. On the left (panels a) and b)), the level curves of $m$ for different values of $\gamma_{\ell}$ are evaluated at the inactive phase. The colored areas correspond to the subspace where the level curves are meaningful. The straight lines correspond to the portions of the plane with fixed non-normality ($\zeta=cont$ or $\xi=cont$). The cyclic and the feedforward setups are respectively addressed in panels a) and b). In both cases, it can be observed that increasing $\gamma_{\ell}$ means shifting the $m$ level curves to the left. Then, for each set ($\gamma_{\nu}$,$\gamma_{\mu}$) $m$ is enhanced. As a final remark, there it appears clear how all the possible values of $m$ can be obtained while still fixing the non-normality parameter.

### v2

@@ III.1.1 Cyclic coupling
In the cyclic architecture ($c=\epsilon$), the system’s symmetry ensures that the equations for each node are identical. Consequently, the equilibrium condition $\bar{x}_{i}=\bar{x}_{j}$ holds for all pairs of indices $i,j$. Moreover, following a reasoning as the one above, one can show that, as before, excitatory and inhibitory activities also reach the same steady state, $\bar{x}_{i}=\bar{y}_{i}$. Therefore, as in the decoupled case, the dynamics can be effectively described by a one-dimensional equation.
The generalization of the decoupled system to the cyclic architecture is straightforward: it is sufficient to replace $\gamma_{\mu}$ with $\gamma_{\mu}+\gamma_{\ell}$. Consequently, the bifurcation point becomes
$\gamma_{\mu}^{c,\epsilon}=\gamma_{\nu}+\alpha-\gamma_{\ell},$ (8)
and the scaling of the non-zero solution is
$\bar{x}\sim\frac{1}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c,\epsilon}\right),$ (9)
i.e., as in the decoupled case, the level of activity grows linearly with the distance to the bifurcation point (thus, the corresponding critical exponent is $\beta=1).$
@@
Fig. 2(a) and (b) show the bifurcation diagram for the population $x_{1}$ and the scaling of the fixed point around the bifurcation point, demonstrating that the curve for the cyclic system as a function of $\gamma_{\mu}$ coincides with that of the decoupled system once appropriately shifted (see figure caption). Moreover, the figure illustrates the scaling of $x_{1}$ as a function of the distance to the bifurcation point, with an exponent $\beta=1$. The figure also illustrates that this scaling exponent holds not only for the cyclic and decoupled systems, but also for a broad range of values $c<\epsilon$, corresponding to architectures that are not perfect cycles, which we refer to as "weak-feedback systems."
These results suggest that cyclic interactions do not fundamentally alter the phase transition of the system, with respect to the uncoupled case, provided that feedback is preserved ($c>0$).
III.1.2 Purely feedforward coupling
In the feedforward system, the units are organized in a chain. We label the populations such that $x_{3}$ is the upstream or initial node, projecting to $x_{2}$, which in turn projects to $x_{1}$. The population $x_{3}$ behaves as a decoupled unit, becoming active at the bifurcation point $\gamma_{\mu}=\gamma_{\mu}^{c}$, and for $\gamma_{\mu}\geq\gamma_{\mu}^{c}$, activation of $x_{3}$ triggers a cascade.
Around the bifurcation point, the dynamics of unit $i$ driven by unit $i+1$ can be expanded as:
$\dot{\text{x}_{i}}$ $\approx\bar{\mathrm{f}}-\alpha\text{x}_{i}-\bar{\mathrm{f}}\text{x}_{i}+\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}-\gamma_{\nu})\text{x}_{i}$ (10)
@@ III.1.2 Purely feedforward coupling
where $\bar{\mathrm{f}}$ and its derivatives are evaluated for the averaged external input $s=\epsilon\gamma_{\ell}\bar{\text{x}}_{i+1}$ (except for $i=3$, that has no input). By defining $\delta\equiv\gamma_{\mu}-\gamma_{\mu}^{c}$, we map this to the normal form $\dot{x}=c+ax^{2}$. The constant driving term scales as $c\propto\bar{\mathrm{f}}\propto\bar{x}_{i+1}$.
Since $x_{3}$ scales linearly with $\delta$ (standard transcritical scaling), $x_{2}$ undergoes a saddle-node bifurcation driven by $x_{3}$, yielding:
$\bar{\text{x}}_{2}\sim\frac{\sqrt{\epsilon\gamma_{\ell}}}{\alpha}\sqrt{\gamma_{\mu}-\gamma_{\mu}^{c}}\sim\delta^{1/2}.$ (11)
Recursively, $x_{1}$ is driven by $x_{2}$, leading to a fourth-root scaling:
$\bar{\text{x}}_{1}\sim\frac{(\epsilon\gamma_{\ell})^{3/4}}{\alpha}\left(\gamma_{\mu}-\gamma_{\mu}^{c}\right)^{1/4}\sim\delta^{1/4}.$ (12)
Generalizing this recursive argument to a feedforward chain of length $N$, the fixed point of the $k$-th node in the chain (where $k=1$ is the input and $k=N$ is the output) must scale as:
$\bar{x}_{k}\sim(\gamma_{\mu}-\gamma_{\mu}^{c})^{1/2^{k-1}},$ (13)
i.e. the exponent $\beta$ depends one the depth of the activated area. This result implies a rapid "flattening" of the response curve as one moves deeper into the chain. For deep feedforward circuits, the downstream nodes become effectively binary —switching sharply between inactive and active states— thereby acting as robust signal rectifiers despite the linear nature of the initial upstream perturbation.
Fig. 2(a/b) show the bifurcation diagram and the scaling of the fixed points in numerical simulations, revealing a perfect agreement with analytical predictions. Comparing the feedforward and cyclic cases, we observe that the fixed point in the feedforward setup is more sensitive to variations in $\gamma_{\mu}$ than in the cyclic architecture.
@@ III.2 Non-normal dynamics in non-reciprocal motifs
Non-normal matrices, i.e. matrices that do not commute with their adjoint (the transpose in the real case), typically have non-orthogonal eigenvectors. In particular, the angle between their right eigenvectors (measured with the standard Euclidean inner product) provides a convenient proxy for non-normality: the smaller the angle, the more pronounced the associated transient effects. Motivated by this, we analyze the eigenvectors of the Jacobian near the equilibrium points and introduce a parameter that quantifies the angle between right eigenvectors. This characterization is key to understanding the phenomenology of the stochastic regime and to highlighting the differences between the cyclic and feedforward architectures.
Let us first rewrite the system in terms of the previously defined variables, $\Sigma$ and $\Delta$
@@
Since the solution is always of the form $\text{x}_{i}=\text{y}_{i}$ $\forall i$, then $\bar{\Delta}_{i}=0$. Therefore, breaking down $\Sigma_{i}$ and $\Delta_{i}$ into the equilibrium and perturbation components, $\Sigma_{i}=\bar{\Sigma}_{i}+\sigma_{i}$ and $\Delta_{i}=\bar{\Delta}_{i}+\delta_{i}$, one can rewrite the linear expansion of the equations as follows:
$\dot{\sigma}_{i}=-(\alpha+\bar{\mathrm{f}})\sigma_{i}+(1-\bar{\Sigma}_{i})\bar{\mathrm{f}}^{\prime}\biggl[\sigma_{i}(\gamma_{\mu}-\gamma_{\nu})+\delta_{i}(\gamma_{\mu}+\gamma_{\nu})+$ (15)
$\hskip 165.02597pt+\epsilon\gamma_{\ell}(\sigma_{i-1}+\delta_{i-1})\biggr]$
$\dot{\delta}_{i}=-(\alpha+\bar{\mathrm{f}})\delta_{i}.$
Note that, in the new coordinate system, the Jacobian associated with the linearized dynamics assumes an upper-triangular form. This change of basis, originally proposed in [5] for symmetry considerations in the decoupled system, effectively implements the Schur transformation of the Jacobian of Eq. (5), guaranteeing its upper-triangular structure in the transformed variables.
The Jacobian matrices are computed for generic fixed points, as our goal is to derive measures of non-normality applicable across regimes.
III.2.1 Cyclic couplings
Let us begin with the study of the cyclic system ($c=\epsilon$). The Jacobian of the cyclic model has a block circulant matrix shape and thus can be block diagonalized according to the circulant matrix theorem [11], as shown in Supp. Mat. A.1. It is interesting to note that the diagonalization of the blocks leads to the basis $(\Sigma_{i},\Delta_{i})$. The same result can be achieved by a Schur transformation in the complex field.
A generic circulant matrix is diagonalized by a special unitary matrix known as the discrete Fourier transform matrix, which has $(r,s)$ element equal to $\exp(-2\pi\mathrm{i}(r-1)(s-1)/n)$. In our model, one can write the diagonalized Jacobian for the cyclic $J_{c}^{(D)}$ as follows:
@@ III.2.1 Cyclic couplings
[…]
$\Psi=\begin{pmatrix}\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}&\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}\gamma_{\ell}.\\ 0&0\end{pmatrix}$.
Therefore, according to this base, the neural circuit can be remapped into three disjoint feedforward graphs as depicted in Fig. 1b. Using the upper triangular shape of the sub-matrices, we can calculate the eigenvalues in the diagonal elements:
@@
where $k_{i}=\biggl\{\epsilon,\epsilon\biggl(-\frac{1}{2}+i\frac{\sqrt{3}}{2}\biggr),\epsilon\biggl(-\frac{1}{2}-i\frac{\sqrt{3}}{2}\biggr)\biggr\}$, $i=1,2,3$. We can identify two complex and four real eigenvalues, wherein global stability is constrained by the inequality $\alpha-\bar{f}+\frac{1}{2}(2-\bar{\Sigma})\bar{\mathrm{f}}^{\prime}(\gamma_{\mu}+\epsilon\gamma_{\ell}-\gamma_{\nu})<0$. In particular the null fixed point loses its stability when $\gamma_{\mu}=\gamma_{\nu}+\alpha$, as already found in Sec. III.1.
Computing the eigenvectors’ elements it becomes evident —-as explicitly shown in in Supp. Mat. A.1.— that $\xi=|{\gamma_{\nu}-\gamma_{\mu}-\epsilon\gamma_{\ell}}|/({\gamma_{\mu}+\gamma_{\nu}+\epsilon\gamma_{\ell}})$ controls the angle between two of the six right eigenvectors. These two eigenvectors encompass contributions of excitatory and inhibitory neurons across all populations; in particular one of them is defined by $\Sigma_{i}=1,\Delta_{i}=0$ –i.e. lies in the direction where all excitatory and inhibitory populations have the same firing rate– and the other one is defined by $\Sigma_{i}=1,\Delta_{i}=\xi$.
Thus, by tuning the parameter $\xi$ one can directly modulate the angle between these two (right) eigenvectors, which determines the degree of non-normality in the system.
In the asymptotic limit $\xi\to\infty$, the eigenvectors become orthogonal and the system approaches normality. Conversely, for $\xi=0$, $J_{c}^{(D)}$ becomes defective, as the two eigenvectors coalesce into one. In this regime, non-normality stems from a single structural mechanism governed solely by $\xi$, which therefore uniquely controls the eigenvector collapse. From this perspective, $\xi$ can be viewed as a generalization of the parameter $\zeta={|\gamma_{\nu}-\gamma_{\mu}|}/({\gamma_{\mu}+\gamma_{\nu}})$, which quantifies non-normality in the decoupled system [12]. Furthermore, the degenerate eigenvector defined by $\Sigma_{i}=1,\Delta_{i}=0$ extends the corresponding structure found in the decoupled system. Once again, by reabsorbing $\epsilon\gamma_{\ell}$ into $\gamma_{\mu}$ the cyclic system maps onto the decoupled one.
Importantly, this result shows that feedback alone does not generate an independent source of structural non-normality: as long as cyclic symmetry is preserved, the collective dynamics remain reducible to those of a single effective E/I unit. This provides a useful baseline against which genuinely non-reducible architectures, such as feedforward chains, can be identified.
III.2.2 Feedforward couplings
We now move on to the feedforward case by setting $c=0$. In this case, the Jacobian matrix takes the form:
@@ III.2.2 Feedforward couplings
where the subscripts indicate that the 22x22 matrices $\Phi_{k}$ (respectively $\Psi_{k}$) share the same structure, but are evaluated at different equilibrium points, since each unit of the chain reaches a different steady state. However, as stated above, the equilibrium point of the excitatory and inhibitory sub-populations within each unit, remains homogeneous (i.e. $\bar{\text{x}}_{i}=\bar{\text{y}}_{i}$). The upper diagonal form of $J_{f}$ allows to read directly the eigenvalues on the diagonal (see Supp. Mat. A.2), i.e.:
@@
Again, the complete form of the eigenvectors, along with their derivation and a generalization to $n$ populations, are provided in Supp. Mat. A.2.
Interestingly, both the elements of the eigenvectors and the eigenvectors themselves exhibit a structured pattern that reflects a recursive-like organization, explicitly highlighting the directionality of signal propagation along the chain. By comparing the pairs of eigenvectors in the Supp. Mat. A.2, two contributions to the system’s non-normality can be identified. The first, encoding the so-called inner non-normality, is given by $\zeta=|{\gamma_{\nu}-\gamma_{\mu}}|/({\gamma_{\mu}+\gamma_{\nu}})$ which coincides with the non-normality parameter of the decoupled system [12]. Similar to the cyclic setup, $J_{f}$ becomes defective for $\zeta\rightarrow 0$ and normal in the limit $\zeta\rightarrow\infty$. Specifically, if $\zeta=0$ the eigenvectors in Supp. Mat. A.2., collapse in pairs. The second contribution, representing the outer non-normality, is associated with the directionality of the flow along the chain, and coincides with the coupling parameter
$\gamma_{\ell}$
. If, in addition to $\zeta=0$, the inter-layer coupling is strong $(\gamma_{\ell}\gg\gamma_{\mu},\gamma_{\nu})$, then all six eigenvectors nearly collapse, and the feedforward architecture yields a dramatic amplification of non-normality. However, if $\zeta\neq 0$, then no pairs of eigenvectors collapse: $J_{f}$ cannot become defective by the effect of the directed network architecture alone.
Figure 3: Reactivity and comparison of numerical abscissa and non-normality parameters. Left panels. (a,b): Level curves of the numerical abscissa $m$ for different values of $\gamma_{\ell}$, evaluated at the inactive phase. Colored regions indicate the subspace where the level curves are defined. Straight lines correspond to fixed values of the non-normality parameters ($\zeta=\text{const}$ or $\xi=\text{const}$). Panel (a) shows the cyclic setup, while panel (b) shows the feedforward setup. Increasing $\gamma_{\ell}$ shifts the $m$ level curves leftward, so that $m$ increases for each pair $(\gamma_{\nu},\gamma_{\mu})$, demonstrating that all possible values of $m$ can be obtained while keeping the non-normality parameter fixed. Right panels (c–f): Time evolution of $\rho(t)$ under variable reactive conditions. Top row (c,d) corresponds to the decoupled system, with couplings fixed and $\xi$ constant. Bottom row (e,f) corresponds to setups where the non-normality of feedforward and cyclic systems is adjusted to maintain a consistent level across configurations. In this case, only the feedforward system exhibits a significant increase in reactivity. Reactivity increases from left to right with the numerical abscissa $m$.